# First Study of the Nuclear Response to Fast Hadrons via Angular Correlations between Pions and Slow Protons in Electron–Nucleus Scattering

### v1

@@ Abstract
We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus ($eA$) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark. The experiment employed the CLAS detector with a 5-GeV electron beam incident on deuterium, carbon, iron, and lead targets. For heavier nuclei, the pion-proton correlation function is more spread-out in azimuth than for lighter ones, and this effect is more pronounced in the $\pi p$ channel than in earlier $\pi\pi$ studies. The proton‑to‑pion yield ratio likewise rises with nuclear mass, although the increase appears to saturate for the heaviest targets. These trends are qualitatively reproduced by state‑of‑the‑art $eA$ event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing. At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in $eA$ scattering.
@@ 1 Introduction
In this study, we examine electron–nucleus ($eA$) DIS events that contain a leading $\pi^{+}$ accompanied by a secondary proton, as illustrated in Fig. 1. Correlating these two hadrons provides new insight into how the nucleus responds to a fast quark or hadron and how the hadronization process is modified within the nucleus. Most of the observed protons come from fragmentation of the nuclear remnant. Because the proton’s rest energy is already present in the initial state, a smaller amount of energy from the hadronic cascade is necessary to knock a proton out of a nucleus than produce a secondary pion with the same momentum. This allows the current measurement to probe lower energy scales of the hadronization process and transport properties of the nucleus than in existing dipion measurements PhysRevC.111.035201.
@@
The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh, since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in $eA$ collisions AbdulKhalek:2021gbh, Chang:2022hkt, RoblesGajardo:2022efe.
@@ 2 Experimental setup
The CLAS CLAS:2003umf detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters (CC), and an electromagnetic calorimeter (EC). Its polar angular acceptance ranged from 8 to 140 for the drift chambers and TOF, and from 8 to 45 for the Cherenkov counters and calorimeter. The drift chambers functioned in a toroidal magnetic field up to 2 Tm, where the magnet’s polarity made negatively charged particles bend toward the beam pipe. The resulting resolutions for charged particles were $\sigma_{p}/p=0.5\%$ for the momentum, 2 mrad for the polar angle, and 4 mrad for the azimuthal angle. The data were selected with a trigger that required at least one electron candidate with momentum $p>500$ MeV/c/c.
@@ 3 Event selection and observables
Following Refs. CLAS:2012tlh, CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC. Background from $\pi^{-}$ was suppressed to the $1\%$ level using the CC and the EC. Charged pions and protons were identified with TOF hits consistent with those types of hadrons with momentum determined in the DCs. The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the CC to suppress proton background. Fiducial cuts on momentum and angles were used in order to avoid regions with steeply varying acceptance or low resolution. The protons’ energies were corrected for energy loss in the target using the same procedure as was used in Ref. CLAS:2022oux.
@@ 5 Results and discussion
We further note that, except when $\Delta Y$ is small and $|\Delta\phi|\approx\pi$, the correlation functions are larger for the nuclear targets than for deuterium (as indicated by the ratio, $R$, being greater than one). Since the correlation functions are related to the proton yield per leading pion, $N_{e^{\prime}\pi p}/N_{\pi p}$, (see Eq. 3), it follows that more protons are produced per leading-$\pi^{+}$ event in nuclei than in deuterium. This can partially be explained by the fact that there are more protons that can be knocked out of a larger nucleus than a smaller nucleus. However, this trend appears to saturate, as the values for the correlation functions for iron and lead are within about 30% of one another despite lead having more than three times as many protons as iron. In a scenario where only the protons in or near the path of the cascade could be knocked out, the number of protons that could be knocked out would scale with the nuclear radius, which can be proxied by $\sqrt[3]{A}$, rather than the atomic number, $Z$. In this case, about ($\sqrt[3]{A_{\rm Pb}}/\sqrt[3]{A_{\rm Fe}}=\sqrt[3]{208/56}\approx 1.55$ times as many protons could be knocked out from Pb than from Fe, which is closer to what we observe than if this were scaled with $Z$. Further, the cascade has a limited amount of energy, some of which causes protons to be knocked out and part of which goes into producing other particles, therefore limiting the number of protons that can be knocked out per reaction. This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert:2020akg, which used D, C, Al, Cu, Sn, and Pb targets. According to this interpretation, the saturation effect should occur at a larger value of $A$ for higher energies than at lower energies.
@@
From the correlation functions, we determined the RMS widths, $\sigma$, integrated in $\Delta Y$, which we show in the top left panel of Fig. 3 as a function of $\sqrt[3]{A}$ (which proxies the radius of the nucleus). We note that $\sigma$ increases monotonically with respect to increasing nuclear size, with values of about 1.13, 1.32, 1.41, and 1.45 for D, C, Fe, and Pb, respectively. This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of $A$, as shown in the lower left panel of the same figure. This trend is similar to what was observed in the dipion measurements PhysRevC.111.035201 (gray open squares in Fig. 3). We note that this increase in width with respect to nuclear size is steeper in this work than in Ref. PhysRevC.111.035201. This seems to suggest that low-energy protons undergo stronger final-state interactions than secondary pions. In the context of some models, such as the GiBUU model Buss:2011mx, when secondary pions are produced, they have a “pre-hadronic” phase when they are color neutral but are not fully formed. During this phase, the “pre-hadrons” have a weaker interaction with the nuclear material than after they have fully formed. Since protons from knockout are already fully formed, they do not have this pre-hadronic phase; therefore they interact with the nucleus as soon as they are struck by the cascade. Another possible explanation is that the protons have lower kinetic energy than pions of the same momentum, and are therefore more susceptible to energy loss and absorption effects.
@@ 6 Summary and conclusions
Compared with earlier dipion measurements PhysRevC.111.035201, the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation. The trends are qualitatively reproduced by the state-of-the-art $eA$ event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.

### v2

@@ Abstract
We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus (eAeA) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark. The experiment employed the CLAS detector with a 5-GeV electron beam incident on deuterium, carbon, iron, and lead targets. For heavier nuclei, the pion-proton correlation function is more spread-out in azimuth than for lighter ones, and this effect is more pronounced in the $\pi p$ channel than in earlier $\pi\pi$ studies. The proton‑to‑pion yield ratio likewise rises with nuclear mass, although the increase appears to saturate for the heaviest targets. These trends are qualitatively reproduced by state‑of‑the‑art eAeA event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing. At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in eAeA scattering.
@@ 1 Introduction
In this study, we examine electron–nucleus (eAeA) DIS events that contain a leading $\pi^{+}$ accompanied by a secondary proton, as illustrated in Fig. 1. Correlating these two hadrons provides new insight into how the nucleus responds to a fast quark or hadron and how the hadronization process is modified within the nucleus. Most of the observed protons come from fragmentation of the nuclear remnant. Because the proton’s rest energy is already present in the initial state, a smaller amount of energy from the hadronic cascade is necessary to knock a proton out of a nucleus than produce a secondary pion with the same momentum. This allows the current measurement to probe lower energy scales of the hadronization process and transport properties of the nucleus than in existing dipion measurements PhysRevC.111.035201.
@@
The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh, since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in eAeA collisions AbdulKhalek:2021gbh, Chang:2022hkt, RoblesGajardo:2022efe.
@@ 2 Experimental setup
The CLAS CLAS:2003umf detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters ($C$), and an electromagnetic calorimeter (EC). Its polar angular acceptance ranged from 8 to 140 for the drift chambers and TOF, and from 8 to 45 for the Cherenkov counters and calorimeter. The drift chambers functioned in a toroidal magnetic field up to 2 Tm, where the magnet’s polarity made negatively charged particles bend toward the beam pipe. The resulting resolutions for charged particles were $\sigma_{p}/p=0.5\%$ for the momentum, 2 mrad for the polar angle, and 4 mrad for the azimuthal angle. The data were selected with a trigger that required at least one electron candidate with momentum $p>500$ MeV/c/c.
@@ 3 Event selection and observables
Following Refs. CLAS:2012tlh, CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC. Background from $\pi^{-}$ was suppressed to the $1\%$ level using the $C$ and the EC. Charged pions and protons were identified with TOF hits consistent with those types of hadrons with momentum determined in the DCs. The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the $C$ to suppress proton background. Fiducial cuts on momentum and angles were used in order to avoid regions with steeply varying acceptance or low resolution. The protons’ energies were corrected for energy loss in the target using the same procedure as was used in Ref. CLAS:2022oux.
@@ 5 Results and discussion
We further note that, except when $\Delta Y$ is small and $|\Delta\phi|\approx\pi$, the correlation functions are larger for the nuclear targets than for deuterium (as indicated by the ratio, $R$, being greater than one). Since the correlation functions are related to the proton yield per leading pion, $N_{e^{\prime}\pi p}/N_{\pi p}$, (see Eq. 3), it follows that more protons are produced per leading-$\pi^{+}$ event in nuclei than in deuterium. This can partially be explained by the fact that there are more protons that can be knocked out of a larger nucleus than a smaller nucleus. However, this trend appears to saturate, as the values for the correlation functions for iron and lead are within about 30% of one another despite lead having more than three times as many protons as iron. In a scenario where only the protons in or near the path of the cascade could be knocked out, the number of protons that could be knocked out would scale with the nuclear radius, which can be proxied by $\sqrt[3]{A}$, rather than the atomic number, $Z$. In this case, about ($\sqrt[3]{A_{\rm Pb}}/\sqrt[3]{A_{\rm Fe}}=\sqrt[3]{208/56}\approx 1.55$ times as many protons could be knocked out from Pb than from Fe, which is closer to what we observe than if this were scaled with $Z$. Further, the cascade has a limited amount of energy, some of which causes protons to be knocked out and part of which goes into producing other particles, therefore limiting the number of protons that can be knocked out per reaction. This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert:2020akg, which used D, C, Al, Cu, Sn, and Pb targets. According to this interpretation, the saturation effect should occur at a larger value of AA for higher energies than at lower energies.
@@
From the correlation functions, we determined the RMS widths, $\sigma$, integrated in $\Delta Y$, which we show in the top left panel of Fig. 3 as a function of $\sqrt[3]{A}$ (which proxies the radius of the nucleus). We note that $\sigma$ increases monotonically with respect to increasing nuclear size, with values of about 1.13, 1.32, 1.41, and 1.45 for D, C, Fe, and Pb, respectively. This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of AA, as shown in the lower left panel of the same figure. This trend is similar to what was observed in the dipion measurements PhysRevC.111.035201 (gray open squares in Fig. 3). We note that this increase in width with respect to nuclear size is steeper in this work than in Ref. PhysRevC.111.035201. This seems to suggest that low-energy protons undergo stronger final-state interactions than secondary pions. In the context of some models, such as the GiBUU model Buss:2011mx, when secondary pions are produced, they have a “pre-hadronic” phase when they are color neutral but are not fully formed. During this phase, the “pre-hadrons” have a weaker interaction with the nuclear material than after they have fully formed. Since protons from knockout are already fully formed, they do not have this pre-hadronic phase; therefore they interact with the nucleus as soon as they are struck by the cascade. Another possible explanation is that the protons have lower kinetic energy than pions of the same momentum, and are therefore more susceptible to energy loss and absorption effects.
@@ 6 Summary and conclusions
Compared with earlier dipion measurements PhysRevC.111.035201, the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation. The trends are qualitatively reproduced by the state-of-the-art eAeA event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.